RELATIVE INTEGRAL CROSS SECTIONS FOR T → R ENERGY TRANSFER IN He–CO COLLISIONS

\underline{ANTONIS P. TSAKOTELLIS}, STILIANA ANTONOVA, AO LIN, AND GEORGE C. MCBANE, Department of Chemistry, The Ohio State University, Columbus, OH 43210.

T → R energy transfer between He and CO has been studied in a crossed beam experiment. CO was initially prepared in its lowest few rotational states by seeded supersonic expansion. Postcollision populations of higher CO rotational states were determined with resonance enhanced multiphoton ionization. Extraction of populations and cross sections from the REMPI signal intensities, comparisons with related experiments, and implications for the He–CO potential surface will be presented.